# Exploiting synergies between JWST and cosmic 21-cm observations to uncover star formation in the early Universe

### v1

@@ 3 Observational data
Models created with 21cmSPACE are often utilized to constraint astrophysical parameters using various observational datasets, as summarized in Table 6 (listed for ease of parsing previous literature). We build upon these works by including the UVLF as an additional observable, and by exploring the effect of the time-evolving enhancement of SFE. We describe all the datasets used in our analysis below.
@@ 3.1 SARAS 3
The Shaped Antenna measurement of the background RAdio Spectrum (SARAS) experiments are a series of radio experiments designed to detect the global 21-cm signal. We use measurements of the global sky temperature from the latest experiment, SARAS 3, in the band 555555\,55 $\,85\,\mathrm{MHz}$ ($z\approx 15\,$ 2525\,2525; Jishnu Nambissan et al., 2021) [^7] We do not include SARAS 2 measurements (Singh et al., 2017) in this work due to uncertainty in the systematics modelling in the dataset (see Bevins et al., 2024, for a joint analysis of both SARAS 2 and 3).. The data consists of 15 hours of observation from a 14-day period during its deployment in Dandiganahalli Lake and Sharavati backwaters in Southern India (Singh et al., 2022). The data is calibrated to remove any radio-frequency interference (RFI), receiver systematics, and thermal emission from the lake. It is thus expected to consist of the global 21-cm signal, the time-averaged foreground from both galactic and extra-galactic sources, and any residual noise.
@@ 3.2 HERA
The Hydrogen Epoch of Reionization Array (HERA) is a radio interferometer with cross-dipole feeds and the PAPER correlator (in its Phase 1), designed to detect the 21-cm power spectrum in the frequency range 100100100\,100 $\,200\,\mathrm{MHz}$ (DeBoer et al., 2017). We use the upper limits on the 21-cm power spectrum from HERA Phase 1 observations (HERA Collaboration et al., 2023), which improve upon the previous limits from HERA Collaboration et al. (2022b). This dataset contains 94 nights of observations using 353535\,35 4141\,4141 antennas across the experiment’s entire frequency range. We use the two frequency bands least contaminated by radio frequency interference, Band 1 in Field D (6.256.256.25\,6.25 9.259.25\,9.259.25 hours LST) and Band 2 in Field C (4.04.04.0\,4.0 6.256.25\,6.256.25 hours LST) in the range 117.19117.19117.19\,117.19 $\,133.11\,\mathrm{MHz}$ and 152.25152.25152.25\,152.25 $\,167.97\,\mathrm{MHz}$ respectively. The lowest limits in the two bands are:
@@
We decimate the data to only include every other $k$-bin to ensure neighbouring data points are uncorrelated in our analysis, as done in Pochinda et al. (2024). As HERA employs a foreground avoidance technique, no foreground modelling is required and we marginalise residual systematics above the thermal noise using the likelihood function described in Section 4.1, as done previously in HERA Collaboration et al. (2023).
@@ 3.3 Cosmic X-ray background
The cosmic X-ray background (CXB) is the unresolved X-ray flux from sources outside of the Milky Way Galaxy (see Bambi & Sangangelo, 2022, for a recent review). CXB includes contribution from point sources such as AGN and galaxies that are too faint to be resolved by telescopes, diffuse emission from the hot IGM, and redshifted emission from X-ray binaries both local and distant (barring unknown instrumental systematics). Hence, measurements of the CXB should provide upper limits for the X-ray emissivity of XRBs in the early Universe and complement the inference on star-formation efficiency (Equation 16).
@@ 3.4 Cosmic radio background
For our analysis, we use the CRB data collated in Table 2 of (Dowell & Taylor, 2018), which includes measurements from LWA1 Low Frequency Sky Survey (Dowell et al., 2017) in the range 404040\,40 $\,80\,\mathrm{MHz}$, ARACDE2 (Fixsen et al., 2011) in the range 333\,3 $\,11\,\mathrm{GHz}$, and other single frequency experiments at $22\,\mathrm{MHz}$ (Roger et al., 1999), $46\,\mathrm{MHz}$ (Alvarez et al., 1997; Maeda et al., 1999), $408\,\mathrm{MHz}$ (Haslam et al., 1982; Remazeilles et al., 2015), and $1420\,\mathrm{MHz}$ (Reich et al., 2001).
Figure 3: UVLF data used for constraints in this work. The data points are from Bouwens et al. (2021a, B21, in purple) and Donnan et al. (2024, D24, in pink), and represent some of the largest compilation of surveys with HST and JWST, respectively. The dashed lines are analytic Schechter and double power law fits from the respective works.
@@ 3.5 UV luminosity function
  • Bouwens et al., 2021a (B21): This dataset combines several HST surveys to provides UVLFs in the range $z=2\,$ 1010\,1010. This includes the HUDF/XDF and parallel fields, BORG/HIPPIES, five CANDELS fields, HFF and parallel fields, and observations from the ERS program (see their Table 1 for details). For our analysis, we use the step-wise UVLF estimates from their Table 4 in the redshift range $z\approx 6\,$ 99\,99 containing 1066, 601, 246 and 33 sources respectively, and omit the 9 sources at $z\approx 10$ from Oesch et al. (2018) in favour of the larger sample size in JWST data at this limiting redshift.
  • Donnan et al., 2024 (D24): This multi-field dataset combines several JWST surveys, including PRIMER, NGDEEP and JADES to provide UVLFs in the redshift range $z\approx 9\,$ 1515\,1515. We pick this dataset due to its inclusion of latest JWST data, its large survey area of $\simeq 370\,\text{arcmin}^{2}$, and its consistent findings with other datasets at both $z=11$ (McLeod et al., 2024; Leung et al., 2023; Pérez-González et al., 2023) and $z=12.5$ (Adams et al., 2024; Robertson et al., 2024). For our analysis, we use UVLF estimates from the entire dataset given in their Table 2, consisting of a total of 2548 sources, in the redshift bins $9.58.5<z<$ 9.5, $10.59.5<z<$ 10.5, $11.510.5<z<$ 11.5, $13.511.5<z<$ 13.5 and a tentative $15.513.5<z<$ 15.5 bin based on an equivalent of $\sim 1.3$ sources888 This is the total contribution from summed probability $p(z)$ of all photometric solutions; for more details on the sample selection process, see Section 3 of their work.. However, note that due to the large error bars of the highest-$z$ data point, it does not affect our constraints strongly.
@@ 4 Bayesian inference
Table 2: Summary of neural network emulator training parameters for each observable: 21-cm global signal, 21-cm power spectrum, cosmic X-ray background, cosmic radio background, and UV luminosity function. All emulators are trained with the Adam optimization algorithm using a $\tanh$ activation function, a mean squared error loss function, and a test/train split of $1/10$. The $N_{\text{samp}}$ parameter denotes the number of sample points per simulation used in the emulator training (e.g. $T_{21}(z)$ consists of 250 redshift points per simulation for the global 21-cm signal emulator). The emulator accuracy quoted here is the maximum $1\sigma$ error in the part of the input space occupied by data (e.g., the SARAS 3 band $z=15-25$ for global 21-cm signal). For a complete error analysis of the emulators, see Appendix B.
@@
21-cm global signal globalemu 16-16-16-16 250 250 $<11\%$
21-cm power spectrum scikit-learn (MLPRegressor) 100-100-100-100 10000 20000 $<13\%$
Cosmic X-ray background globalemu 16-16-16-16 400 400 $<5\%$
Cosmic radio background globalemu 16-16-16-16 100 100 $<4\%$
UV luminosity function scikit-learn (MLPRegressor) 100-100-100-100 12000 15000 $<2\%$
@@
[…]
@@
[…]
@@ 4.1 Emulator training
We train emulators for our five observables of interest: the 21-cm global signal $T_{21}(z)$, the 21-cm power spectrum $\Delta_{21}^{2}(k,z)$, the cosmic X-ray background $S_{X}(E)$, cosmic radio background $T_{\text{r}}(\nu)$ and UV luminosity function $\Phi(M_{\text{UV}},z)$. We use the Tensorflow (Abadi et al., 2016) based package globalemu (Bevins et al., 2021) to train the one-dimensional observables: 21-cm global signal, CRB and CXB, and scikit-learn to train the two-dimensional observables: 21-cm power spectrum and UVLF. All emulators are trained with the Adam optimization algorithm using a $\tanh$ activation function, a mean squared error loss function, and a test/train split of $1/10$. The input pre-processing and training details for each emulator are as follows:
@@
  • 21-cm global signal: We train the emulator using 21-cm global signal $T_{21}$ output from the simulations in the range $z=6\,$ 3030\,3030. We pre-process the training data by first resampling the default simulation redshift step of $\Delta z=1$ to $\Delta z=0.1$ ($N_{\text{samp}}=250$) via simple interpolation to increase the training dataset size and resolution. We then down-scale and normalize the signals (subtracting the mean and dividing by the standard deviation of training data), as is a standard feature in globalemu. The emulator is trained on a 4-layer NN of 16 nodes each, a batch-size of 250, and with early-stopping based on validation loss to prevent over-fitting. The resulting emulator has an accuracy of $<11\%$ error in the SARAS 3 band at 68% confidence, and an RMSE of $\sim 70\,\text{mK}$ across the entire redshift range at the 95th percentile of test data. For a more detailed analysis of emulator accuracy (and its dependence on the input-space $z$), see Appendix B.
@@
  • 21-cm power spectrum: We use a similar methodology here as described in Pochinda et al. (2024) using scikit-learn. The simulation outputs the 21-cm power spectrum $\Delta_{21}^{2}$ at redshifts $z=7\,$ $1$\,$1$ and wavevector range $k\approx 0.1\,$ $\,2.5\,h\,\mathrm{Mpc}^{-1}$. We pre-process the training data by log-transforming both $z$ and $k$, as well as $\Delta_{21}^{2}$ (truncating values $<10^{-8}\,\text{mK}^{2}$) to improve the performance of the emulator over the large dynamic range. We then resample and interpolate the power spectrum to a finer input grid by drawing $N_{\text{samp}}=20000$ samples from uniform pairs of $z$ and $k$ values. The emulator is trained using an MLPRegressor with 4 hidden layers of 100 nodes each and a batch-size of 10000. The resulting emulator has an accuracy of $<13\%$ in linear space in the $k$ and $z$ regime of HERA data at 68% confidence.
@@
  • CXB / CRB: We train the cosmic X-ray background $S_{X}$ and radio background $T_{\text{r}}$ in a similar manner to the 21-cm global signal. In case of the former, the CXB (Equation 18) is output by the simulation in the energy range $E=10^{-1}\,$ $\,10^{3}\,\mathrm{keV}$ with $N_{\text{samp}}=400$. The CRB on the other hand is calculated in post-processing from star-formation rates output by the simulation (Equation 20) in $\nu=10\,$ $\,10^{3}\,\mathrm{MHz}$ with $N_{\text{samp}}=400$ to cover the CRB data comfortably. The data is first log-transformed in both cases, and then normalized as described above for the 21-cm global signal emulator. The emulator training parameters remain the same, except for a batch size of 400 for the CXB emulator and 100 for the CRB emulator. The resulting emulators have an accuracy of $<5\%$ and $4\%$ error in the observational energy bands and frequencies, respectively, in linear space at 68% confidence.
[…]
@@ 4.2 Likelihood evaluation
model $\text{model error}=\sqrt{\text{theory error}^{2}+\text{emulator error}^{2}}.$ (30)
The former accounts for the assumed error on the semi-numerical model used in the simulation, fixed to an ad-hoc 20% (e.g. Zahn et al., 2011), while the latter is derived by the performance of the emulator on test data. For a target function $f(x)$, the emulator performance may vary as a function of $x$. We evaluate this for all emulators in Appendix B, which then feeds into our likelihood functions described below.
@@
    where function $f_{\text{N}}$ normalizes the log-frequency to be in the range $[-1,1]$. Next, we use the emulator to predict the 21-cm global signal at the frequencies of the SARAS 3 data and evaluate the likelihood function:
@@
    where $j$ indexes the observational $k$ and $z$ bins, $\Delta_{\rm 21,HERA}^{2}$ is the observed HERA upper limits, $\Delta_{21,\text{model}}^{2}$ is the theory model predicted 21-cm power spectrum, $\sigma_{\text{obs}}$ is the observational error, and $\sigma_{\text{model}}$ is the model error at those $k_{j}$ and $z_{j}$ values.
  • CXB / CRB: For the CXB and CRB inference, we use a similar likelihood to HERA since we use the data in the form of upper limits. In case of the former, the observable is the CXB flux $S_{X}(E_{\text{band}})$ and we use the emulator to predict this at a high resolution within each energy band of the CXB data and then integrate the flux over the bin. In case of CRB, the observable is the temperature $T_{\text{r}}(\nu)$, and we use the emulator to predict the CRB flux at the precise frequencies of the CRB data. The likelihood function is then evaluated as in Equation 33 for both datasets.
@@
where $\sigma_{+/-}=\sqrt{\left(\pm\sigma_{\text{obs},i,j}\right)^{2}+\left(\sigma_{\text{model}}(M_{\text{UV},i},z_{j})\right)^{2}}$ is the asymmetric error on the UVLF data point, and the denominator of the exponential is $\sigma_{+}$ if $\log_{10}\left(\Phi_{\text{th}}\right)>\log_{10}\left(\Phi_{\text{obs}}\right)$, and $\sigma_{-}$ otherwise.

### v2

@@ 3 Observational data
Models created with 21cmSPACE are often utilized to constrain astrophysical parameters using various observational datasets, as summarized in Table C1 (listed for ease of parsing previous literature). We build upon these works by including the UVLF as an additional observable, and by exploring the effect of the time-evolving enhancement of SFE. We describe all the datasets used in our analysis below.
@@ 3.1 SARAS 3
The Shaped Antenna measurement of the background RAdio Spectrum (SARAS) experiments are a series of radio experiments designed to detect the global 21-cm signal. We use measurements of the global sky temperature from the latest experiment, SARAS 3, in the band 5555\, $85\,\mathrm{MHz}$ ($z\approx 15\,$ 25\,25; Jishnu Nambissan et al., 2021) [^9] We do not include SARAS 2 measurements (Singh et al., 2017) in this work due to uncertainty in the systematics modelling in the dataset (see Bevins et al., 2024, for a joint analysis of both SARAS 2 and 3).. The data consists of 15 hours of observation from a 14-day period during its deployment in Dandiganahalli Lake and Sharavati backwaters in Southern India (Singh et al., 2022). The data is calibrated to remove any radio-frequency interference (RFI), receiver systematics, and thermal emission from the lake. It is thus expected to consist of the global 21-cm signal, the time-averaged foreground from both Galactic and extragalactic sources, and any residual noise.
@@ 3.2 HERA
The Hydrogen Epoch of Reionization Array (HERA) is a radio interferometer with cross-dipole feeds and the PAPER correlator (in its Phase 1), designed to detect the 21-cm power spectrum in the frequency range 100100\, $200\,\mathrm{MHz}$ (DeBoer et al., 2017). We use the upper limits on the 21-cm power spectrum from HERA Phase 1 observations (HERA Collaboration et al., 2023), which improve upon the previous limits from HERA Collaboration et al. (2022b). This dataset contains 94 nights of observations using 3535\, 41\,41 antennas across the experiment’s entire frequency range. We use the two frequency bands least contaminated by radio frequency interference, Band 1 in Field D (6.256.25\, 9.25\,9.25 hours LST) and Band 2 in Field C (4.04.0\, 6.25\,6.25 hours LST) in the range 117.19117.19\, $133.11\,\mathrm{MHz}$ and 152.25152.25\, $167.97\,\mathrm{MHz}$ respectively. The lowest limits in the two bands are:
@@
We reduce the data to only include every other $k$-bin to ensure neighbouring data points are uncorrelated in our analysis, as done in Pochinda et al. (2024). As HERA employs a foreground avoidance technique, no foreground modelling is required and we marginalise residual systematics above the thermal noise using the likelihood function described in Section 4.1, as done previously in HERA Collaboration et al. (2023).
@@ 3.3 Cosmic X-ray background
The cosmic X-ray background (CXB) is the unresolved X-ray flux from sources outside of the Milky Way Galaxy (see Bambi & Sangangelo, 2022, for a recent review). CXB includes contribution from point sources such as AGN and galaxies that are too faint to be resolved by telescopes, diffuse emission from the hot IGM, and redshifted emission from X-ray binaries both local and distant (barring unknown instrumental systematics). Hence, measurements of the CXB should provide upper limits for the X-ray emissivity of XRBs in the early Universe (e.g. Fialkov et al., 2017) and complement the inference on star-formation efficiency (Equation 16).
@@ 3.4 Cosmic radio background
For our analysis, we use the CRB data collated in Table 2 of (Dowell & Taylor, 2018), which includes measurements from LWA1 Low Frequency Sky Survey (Dowell et al., 2017) in the range 4040\, $80\,\mathrm{MHz}$, ARACDE2 (Fixsen et al., 2011) in the range 33\, $11\,\mathrm{GHz}$, and other single frequency experiments at $22\,\mathrm{MHz}$ (Roger et al., 1999), $46\,\mathrm{MHz}$ (Alvarez et al., 1997; Maeda et al., 1999), $408\,\mathrm{MHz}$ (Haslam et al., 1982; Remazeilles et al., 2015), and $1420\,\mathrm{MHz}$ (Reich et al., 2001).
Figure 3: UVLF data used for constraints in this work. The data points are from Bouwens et al. (2021a, B21, in purple) and Donnan et al. (2024, D24, in pink), and represent some of the largest compilations of galaxies with HST and JWST, respectively. The dashed lines are analytic Schechter and double power law fits from the respective works.
@@ 3.5 UV luminosity function
  • Bouwens et al., 2021a (B21): This dataset combines several HST surveys to provides UVLFs in the range z=2z=2\, 10\,10. This includes the HUDF/XDF and parallel fields, BORG/HIPPIES, five CANDELS fields, HFF and parallel fields, and observations from the ERS program (see their Table 1 for details). For our analysis, we use the step-wise UVLF estimates from their Table 4 in the redshift range $z\approx 6\,$ 9\,9 containing 1066, 601, 246 and 33 sources respectively, and omit the 9 sources at $z\approx 10$ from Oesch et al. (2018) in favour of the larger sample size in JWST data at this limiting redshift.
  • Donnan et al., 2024 (D24): This multi-field dataset combines several JWST surveys, including PRIMER, NGDEEP and JADES to provide UVLFs in the redshift range $z\approx 9\,$ 15\,15. We pick this dataset due to its inclusion of the latest JWST data, its large survey area of $\simeq 370\,\text{arcmin}^{2}$, and its consistent findings with other datasets at both $z=11$ (McLeod et al., 2024; Leung et al., 2023; Pérez-González et al., 2023) and $z=12.5$ (Adams et al., 2024; Robertson et al., 2024). For our analysis, we use UVLF estimates from the entire dataset given in their Table 2, consisting of a total of 2548 sources, in the redshift bins $8.5<z<9.5$, $9.5<z<10.5$, $10.5<z<11.5$, $11.5<z<13.5$ and a tentative $13.5<z<15.5$ bin based on an equivalent of $\sim 1.3$ sources101010 This is the total contribution from summed probability $p(z)$ of all photometric solutions; for more details on the sample selection process, see Section 3 of their work.. However, note that due to the large error bars of the highest-$z$ data point, it does not affect our constraints strongly.
@@ 4 Bayesian inference
Table 2: Summary of neural network emulator training parameters for each observable: 21-cm global signal, 21-cm power spectrum, cosmic X-ray background, cosmic radio background, and UV luminosity function. All emulators are trained with the Adam optimization algorithm using a $\tanh$ activation function, a mean squared error loss function, and a test/train split of $1/9$. The $N_{\text{samp}}$ parameter denotes the number of sample points per simulation used in the emulator training (e.g. $T_{21}(z)$ consists of 250 redshift points per simulation for the global 21-cm signal emulator). The emulator accuracy quoted here is the maximum $1\sigma$ error in the part of the input space occupied by data (e.g., the SARAS 3 band z=1525z=15-25 for global 21-cm signal). For a complete error analysis of the emulators, see Appendix B.
@@
21-cm global signal globalemu 16-16-16-16 250 250 $11\%$
21-cm power spectrum scikit-learn (MLPRegressor) 100-100-100-100 10000 20000 $13\%$
Cosmic X-ray background globalemu 16-16-16-16 400 400 $5\%$
Cosmic radio background globalemu 16-16-16-16 100 100 $4\%$
UV luminosity function scikit-learn (MLPRegressor) 100-100-100-100 20000 15000 $2\%$
@@
where $\mathcal{Z}=P(\mathcal{D})=\int\mathcal{L}(\theta)\pi(\theta)d\theta$ is the Bayesian evidence (the integral of the likelihood over the parameter space). A uniform prior is uninformed and suitable for parameters which have an expected range, while a log-uniform prior works better for parameters whose order of magnitude is not known.
@@
$\mathcal{P}(\theta_{i})=\dfrac{1}{\mathcal{Z}}\int\mathcal{L}(\theta)\pi(\theta)d\theta_{j\neq i}.$ (29)
@@ 4.1 Emulator training
We train emulators for our five observables of interest: the 21-cm global signal $T_{21}(z)$, the 21-cm power spectrum $\Delta_{21}^{2}(k,z)$, the cosmic X-ray background $S_{X}(E)$, cosmic radio background $T_{r}(\nu)$ and UV luminosity function $\Phi(M_{\text{UV}},z)$. We use the Tensorflow (Abadi et al., 2016) based package globalemu (Bevins et al., 2021) to train the one-dimensional observables: 21-cm global signal, CRB and CXB, and scikit-learn (Pedregosa et al., 2011) to train the two-dimensional observables: 21-cm power spectrum and UVLF. All emulators are trained with the Adam optimization algorithm using a $\tanh$ activation function, a mean squared error loss function, and a test/train split of $1/9$. The input pre-processing and training details for each emulator are as follows:
@@
  • 21-cm global signal: We train the emulator on the 21-cm global signal $T_{21}$ output from the simulations in the range z=6z=6\, 30\,30. We pre-process the training data by first resampling the default simulation redshift step of $\Delta z=1$ to $\Delta z=0.1$ ($N_{\text{samp}}=250$) via simple interpolation to increase the training dataset size and resolution. We then down-scale and normalize the signals (subtracting the mean and dividing by the standard deviation of training data), as is a standard feature in globalemu. The emulator is trained on a 4-layer NN of 16 nodes each, a batch-size of 250, and with early-stopping based on validation loss to prevent over-fitting. The resulting emulator has an accuracy of $11\%$ error in the SARAS 3 band at 68% confidence, and an RMSE of $\sim 70\,\text{mK}$ across the entire redshift range at the 95th percentile of test data. For a more detailed analysis of emulator accuracy (and its dependence on the input-space $z$), see Appendix B.
@@
  • 21-cm power spectrum: We use a similar methodology here as described in Pochinda et al. (2024) using scikit-learn. The simulation outputs the 21-cm power spectrum $\Delta_{21}^{2}$ at redshifts z=7z=7\, 11\,11 and wavevector range $k\approx 0.1\,$ $2.5\,h\,\mathrm{Mpc}^{-1}$. We pre-process the training data by log-transforming both $z$ and $k$, as well as $\Delta_{21}^{2}$ (truncating values $10^{-8}\,\text{mK}^{2}$) to improve the performance of the emulator over the large dynamic range. We then resample and interpolate the power spectrum to a finer input grid by drawing $N_{\text{samp}}=20000$ samples from uniform pairs of $z$ and $k$ values. The emulator is trained using an MLPRegressor with 4 hidden layers of 100 nodes each and a batch-size of 10000. The resulting emulator has an accuracy of $13\%$ in linear space in the $k$ and $z$ regime of HERA data at 68% confidence.
@@
  • CXB / CRB: We train the cosmic X-ray background $S_{X}$ and radio background $T_{r}$ in a similar manner to the 21-cm global signal. In case of the former, the CXB (Equation 18) is output by the simulation in the energy range $E=10^{-1}\,$ $10^{3}\,\mathrm{keV}$ with $N_{\text{samp}}=400$. The CRB on the other hand is calculated in post-processing from star-formation rates output by the simulation (Equation 20) in $\nu=10\,$ $10^{3}\,\mathrm{MHz}$ with $N_{\text{samp}}=400$ to cover the CRB data comfortably. The data is first log-transformed in both cases, and then normalized as described above for the 21-cm global signal emulator. The emulator training parameters remain the same, except for a batch size of 400 for the CXB emulator and 100 for the CRB emulator. The resulting emulators have an accuracy of $5\%$ and $4\%$ error in the observational energy bands and frequencies, respectively, in linear space at 68% confidence.
  • UVLF: We train the UV luminosity function $\Phi(M_{\text{UV}},z)$ output from the simulation in a similar manner to the 21-cm power spectrum described above. The simulation code outputs the UVLF (in log space) at integer redshifts in the range z=6z=6\, 16\,16 and magnitude $M_{\text{UV}}=\left[-23,-13\right]\,\mathrm{mag}$ with a bin size of $\Delta M_{\text{UV}}=0.5$ to match most observations. The training data is preprocessed by drawing $N_{\text{samp}}=15000$ from uniform pairs of $M_{\text{UV}}$ and $z$ values, and the emulator training follows that of the 21-cm power spectrum with a batch size of $20000$. The resulting emulator has an accuracy of $2\%$ in log space at the observed median $z$ and magnitude bins $M_{\text{UV}}$ at $68\%$ confidence, corresponding to an RMSE of $0.8\,\text{dex}$ across the full magnitude range at 95th percentile of the test data.
@@ 4.2 Likelihood evaluation
$\text{model error}=\sqrt{\text{theory error}^{2}+\text{emulator error}^{2}}.$ (30)
The former accounts for the assumed error on the semi-numerical model used in the simulation, fixed to an ad-hoc 20% (e.g. Zahn et al., 2011; Hutter, 2018), while the latter is derived by the performance of the emulator on test data. For a target function $f(x)$, the emulator performance may vary as a function of $x$. We evaluate this for all emulators in Appendix B, which then feeds into our likelihood functions described below.
@@
    where function $f_{\text{N}}$ normalizes the log-frequency to be in the range [1,1][-1,1]. Next, we use the emulator to predict the 21-cm global signal at the frequencies of the SARAS 3 data and evaluate the likelihood function:
@@
    where $j$ indexes the observational $k$ and $z$ bins, $\Delta_{\rm 21,HERA}^{2}$ is the observed HERA upper limits, $\Delta_{21,\text{model}}^{2}$ is the theory model predicted 21-cm power spectrum, $\sigma_{\text{HERA}}$ is the observational error, and $\sigma_{\text{model}}$ is the model error at those $k_{j}$ and $z_{j}$ values.
  • CXB / CRB: For the CXB and CRB inference, we use a similar likelihood to HERA since we use the data in the form of upper limits. In case of the former, the observable is the CXB flux $S_{X}(E_{\text{band}})$ and we use the emulator to predict this at a high resolution within each energy band of the CXB data and then integrate the flux over the bin. In case of CRB, the observable is the temperature $T_{r}(\nu)$, and we use the emulator to predict the CRB flux at the precise frequencies of the CRB data. The likelihood function is then evaluated as in Equation 33 for both datasets.
@@
where $\sigma_{+/-}=\sqrt{\left(\sigma_{+/-,\text{obs},i,j}\right)^{2}+\left(\sigma_{\text{model}}(M_{\text{UV},i},z_{j})\right)^{2}}$ is the asymmetric error on the UVLF data point, and the denominator of the exponential is $\sigma_{+}$ if $\log_{10}\left(\Phi_{\text{th}}\right)>\log_{10}\left(\Phi_{\text{obs}}\right)$, and $\sigma_{-}$ otherwise.
@@
Figure 4: Marginal 1D and 2D posteriors of astrophysical parameters from the joint analysis and individual analysis of the datasets: UVLF, CXB, CRB, HERA, and SARAS 3 (color-coded as shown in the legend). The 2D posterior contours show the 68% (dash-dotted) and 95% (dashed) credible regions emerging from the joint analysis. The 1D marginal posterior PDFs show the weighted mean (solid line) and 68% credible region (shaded with dash-dotted outline; calculated as the smallest CDF intervals containing the required probability volume) for the joint fit (black), as well as the 1D PDFs for each individual observable (indicated in the legend). Top right panel shows the posterior to prior volume contraction for each dataset, calculated using margarine. For a similar plot comparing joint constraints with and without the UVLF dataset, see Figure 2, and for joint constraints with and without the SARAS 3 dataset, see Figure 3. For the full parameter space, including SARAS 3 nuisance parameters, see Figure 4.
Table 3: Marginal 1D constraints on the parameter $V_{c}$, and related quantities $M_{\text{vir}}$ (Equation 3) and $T_{\text{vir}}$ (Equation 6). The quoted errors are at 68% (95%) confidence levels around the weighted mean. We list SARAS 3 and UVLF individually since they are the most constraining datasets, providing lower and upper limits respectively. The joint analysis combines the two, with SARAS 3 weakening the upper bound on $V_{c}$ set by UVLFs as seen in Figure 3 (or providing a lower bound depending on the foreground fit, as in Figure 6).
Parameter Prior SARAS 3 UVLF Joint w/o SARAS 3 Joint
$\log_{10}\left(V_{c}/\,\mathrm{km}\,\mathrm{s}^{-1}\right)$ $\mathcal{U}\left(0.62,2\right)$ or $\gtrsim 1.16\,(0.72)$ $\lesssim 1.41\,(1.79)$ $\lesssim 1.33\,(1.78)$ $\lesssim 1.40\,(1.79)$
$\implies V_{c}\ (\,\mathrm{km}\,\mathrm{s}^{-1})$ $\log\mathcal{U}(4.2,100)$ $\gtrsim 14.4\,(5.3)$ $\lesssim 25.7\,(62.1)$ $\lesssim 21.3\,(60.4)$ $\lesssim 25.2\,(62.0)$
$\log_{10}\left(M_{\text{vir}}(z=6)/\,\mathrm{M}_{\odot}\right)$ $\mathcal{U}\left(6.5,10.6\right)$ $\gtrsim 8.09\,(6.77)$ $\lesssim 8.84\,(9.99)$ $\lesssim 8.60\,(9.95)$ $\lesssim 8.82\,(9.99)$
$\log_{10}\left(T_{\text{vir}}/\,\text{K}\right)$ $\mathcal{U}\left(2.8,5.6\right)$ $\gtrsim 3.87\,(3.00)$ $\lesssim 4.38\,(5.14)$ $\lesssim 4.21\,(5.12)$ $\lesssim 4.36\,(5.14)$